# MGS-SLAM: Monocular Sparse Tracking and Gaussian Mapping with Depth Smooth Regularization

### v1

@@ Abstract
This letter introduces a novel framework for dense Visual Simultaneous Localization and Mapping (VSLAM) based on Gaussian Splatting. Recently Gaussian Splatting-based SLAM has yielded promising results, but rely on RGB-D input and is weak in tracking. To address these limitations, we uniquely integrates advanced sparse visual odometry with a dense Gaussian Splatting scene representation for the first time, thereby eliminating the dependency on depth maps typical of Gaussian Splatting-based SLAM systems and enhancing tracking robustness. Here, the sparse visual odometry tracks camera poses in RGB stream, while Gaussian Splatting handles map reconstruction. These components are interconnected through a Multi-View Stereo (MVS) depth estimation network. And we propose a depth smooth loss to reduce the negative effect of estimated depth maps. Furthermore, the consistency in scale between the sparse visual odometry and the dense Gaussian map is preserved by Sparse-Dense Adjustment Ring (SDAR). We have evaluated our system across various synthetic and real-world datasets. The accuracy of our pose estimation surpasses existing methods and achieves state-of-the-art performance. Additionally, it outperforms previous monocular methods in terms of novel view synthesis fidelity, matching the results of neural SLAM systems that utilize RGB-D input.
@@ I INTRODUCTION
Simultaneous Localization and Mapping (SLAM) is a key technology in robotics and autonomous driving. It aims to solve the problem of how robots determine their location and reconstruct maps of the environment in unknown scenes. The development of SLAM technology has gone through multiple stages, starting with the initial filter-based methods [1], [2], advancing to graph optimization-based methods[3], [4], and more recently, integrating deep learning. This integration has significantly improved the accuracy and robustness of SLAM systems. With the rapid development of deep learning technology, a new approach to SLAM technology has emerged, utilizing differentiable rendering. The initial applications of differentiable rendering-based SLAM utilized Neural Radiance Fields (NeRF) as their foundational construction method. NeRF, as detailed in [5], employs neural networks to represent 3D scenes, enabling the synthesis of high-quality images and the recovery of dense geometric structures from multiple views. NeRF-based SLAM systems preserve detailed scene information during mapping, which enhances support for subsequent navigation and path planning. However, NeRF’s approach requires multiple forward predictions for each pixel during image rendering, leading to significant computational redundancy. Consequently, this inefficiency prevents NeRF-based SLAM from operating in real-time, thus limiting its practicality for immediate downstream tasks.
@@
Recently, a novel scene representation framework called 3D Gaussian Splatting [6] has demonstrated superior performance compared to NeRF. It features a more concise scene representation method and real-time rendering capability. 3D Gaussian Splatting represents the geometries in a scene using a smooth and continuously differentiable method. This representation method not only delivers an accurate description of the scene but also offers a differentiable approach for optimizing the scene and camera poses. This opens up a new research direction for differentiable rendering-based SLAM. However, current Gaussian Splatting-based SLAM systems rely on the depth maps input from RGB-D cameras, which constrains the scope of their application.
Figure 2: System pipeline. The system inputs an RGB stream and operates frontend and backend processes in parallel. In the frontend, sparse visual odometry extracts patch features from images to estimate poses. These estimated poses and images are inputs to a pre-trained Multi-View Stereo (MVS) network, which estimates priori depth maps. In the backend, the estimated priori depth maps and images, coupled with poses from the frontend, are utilized as supervisory information to construct a Gaussian map. The frontend and backend maintain scale consistency through the Sparse-Dense Adjustment Ring (SDAR) strategy.
This letter presents MGS-SLAM, a novel monocular Gaussian Splatting-based SLAM system. This work introduces several groundbreaking advancements in the field of SLAM, which include integrating Gaussian Splatting-based techniques with sparse visual odometry, employing a pre-trained Multi-View Stereo (MVS) depth estimation network, pioneering a geometric smooth depth loss, and developing the Sparse-Dense Adjustment Ring (SDAR) to ensure scale consistency. Together, these innovations significantly improve the accuracy and functionality of SLAM systems that rely solely on RGB images input. Fig. 1 illustrates the map construction process: initially, sparse visual odometry constructs the sparse map; subsequently, the MVS depth estimation network generates priori depth maps; this depth maps, along with the sparse point map, is then refined through depth optimization in the SDAR; and finally, the Gaussian map is constructed using the optimized priori depth maps and depth smooth regularization loss.
@@
  • Introducing the first SLAM system that combines sparse visual odometry with 3D Gaussian Splatting to achieve the construction of a dense Gaussian map using only RGB images as input.
  • Developing a pre-trained Multi-View Stereo (MVS) depth estimation network that utilizes sparse odometry keyframes and their poses to estimate prior depth maps, thus providing crucial geometric constraints for Gaussian map reconstruction.
@@ II RELATED WORKS
Monocular Dense SLAM. Over the past few decades, monocular dense SLAM technology has seen significant advancements. DTAM [7] pioneered one of the earliest real-time dense SLAM systems by performing parallel depth computations on GPU. To balance computational costs and accuracy, there are also semi-dense methods such as [8], [9], but these methods struggle to capture areas with poor texture. In the era of deep learning, DROID-SLAM [10] utilizes optical flow networks to establish dense pixel correspondences and achieve precise pose estimation. Another study [11], [12] combines a real-time VO/SLAM system with a Multi-View Stereo (MVS) network for parallel tracking and dense depth estimation, and then the Truncated Signed Distance Function (TSDF) [13] is used to fuse depth maps and extract mesh. [14] and [15] incorporate sparse point cloud correction and volumetric fusion strategy on the estimated depth map to mitigate the impact of errors in the estimated depth map. We have also adopted a strategy for correcting the estimated depth map, but the difference is that we use a linear variance correction as depth optimization, which is less computations.
@@
Differentiable Rendering SLAM. With the emergence of Neural Radiance Fields (NeRF) in 2020, numerous NeRF-based SLAM works have been proposed. iNerf [16] was the first to propose optimizing camera poses within a fixed NeRF reconstruction scene, which initiated the SLAM adaptation of NeRF. iMAP [17] represented the pioneering work in NeRF-based SLAM, utilizing a dual-threading mode to track camera poses and execute mapping simultaneously. NICE-SLAM [18] introduced feature grids based on iMAP, enabling NeRF-based SLAM to represent larger scenes. Subsequent works such as GO-SLAM [19], HI-SLAM [20], and Loopy-SLAM [21] incorporated global bundle adjustment (BA) and loop closure correction, further enhancing pose estimation accuracy and mapping performance. Recently, 3D Gaussian Splatting has shown superior performance in 3D scene representation, with its fast rendering capability and Gaussian scene representation being more suitable for online systems like SLAM. SplaTAM [22] combines 3D Gaussian Splatting with SLAM, leveraging the realistic scene reconstruction ability of 3D Gaussian Splatting to surpass NeRF-based SLAM methods in rendering quality. MonoGS [23] achieves monocular map reconstruction of Gaussian Splatting-based SLAM in small scenes through a map clipping strategy. However, existing Gaussian Splatting-based SLAM implementations typically require depth maps input from RGB-D sensors, and their monocular methods are often unable to operate effectively on larger datasets, such as Replica [24]. We propose the system utilizing only RGB images as input and demonstrated enhanced robustness by successfully operating on the Replica dataset.
@@
Figure 3: The priori depth estimation network. The inputs to the network is images with poses from sparse visual odometry, image features are extracted by Feature Pyramid Network (FPN) and warping to 2D cost volume. Finally, encoded and decoded to depth maps using coarse-to-fine strategy.
@@ III METHODS
Our approach utilizes RGB images as input, parallelly performing camera poses estimation and photorealistic dense mapping. As depicted in Fig. 2, the core idea of the approach is to use a pre-trained Multi-View Stereo (MVS) network to couple monocular sparse visual odometry and dense Gaussian Splatting mapping and to maintain scale consistency between sparse point cloud map and dense Gaussian map by the Sparse-Dense Adjustment Ring (SDAR). Specifically, in the frontend part, tracking RGB images provides the backend with coarse camera poses and priori depth maps (Sec. III-A). In the backend part, represent the dense map using Gaussians, and jointly optimize the map and the coarse poses from the frontend (Sec. III-B). In the system components part, system initialization, selecting the keyframes for the system and correcting the scale between sparse point cloud map and dense Gaussian map by SDAR strategy are reported (Sec. III-C).
@@ III-A Sparse Visual Odometry Frontend
To achieve more accurate camera poses tracking and provide dense depth geometry before backend mapping, the frontend of our framework is built on the Deep Patch Visual Odometry (DPVO) [25] algorithm. DPVO is a learning-based sparse monocular visual odometry method for estimating camera poses and building sparse point cloud map by constructing and optimizing errors on patch graph. Given an input RGB stream, the scene is represented as a collection of camera poses $\textbf{T}\in SE(3)^{N}$ and a series of square image patches P extracted from the images, and add them to the bi-directional patch graph that connects the patches to the frames. The reprojection of a square patch $k$ taken from frame $i$ in frame $j$ can be formulated as:
@@
The core of DPVO is an update operator that computes the hidden state for each edge $(k,i,j)\in\varepsilon$. It optimize the reprojection errors on the patch graph to predict a 2D correction vector $\delta^{ij}_{k}\in\mathbb{R}^{2}$ and confidence weight $\psi^{ij}_{k}\in\mathbb{R}^{2}$. Bundle Adjustment (BA) is performed using optical flow correction as a constraint, with iterative updates and refinement of camera poses and patch depths achieved through the non-linear least squares method. The cost function for bundle adjustment is as follows:
@@
Multi-view priori depth estimation. The backend dense Gaussian mapping requires the supervision of geometric depth maps. Unlike the previous method [26], we use a pre-trained Multi-View Stereo (MVS) network to estimate priori depth maps on the keyframes window of DPVO, the network is shown in Fig. 3. Inspired by the monocular depth estimation method [27], our MVS network consists entirely of 2D convolutions with a coarse-to-fine structure that progressively refines the estimated priori depth map to reduce the runtime of the MVS network.
To be more specific, the frame currently tracked by the sparse visual odometry is used as the reference image $\textbf{I}^{0}$. Additionally, we employ the previous N keyframes as a series of original images $\textbf{I}^{n\in{1,...,N}}$. These images and their corresponding camera poses, serve as inputs to the MVS network. Utilizing the Feature Pyramid Network (FPN) module, we extract three layers of image features $\textbf{F}^{s}_{i}$ for each image, with $s$ denoting the layer index and $i$ representing the image index. Each layer contains image features with twice the length and width of the previous layer. In each layer, the image features of the original images dot the image feature of the reference image by a differentiable warping operation to obtain a cost volume with dimensions $\textbf{D}\times\textbf{H}^{s}\times\textbf{W}^{s}$, and the priori depth map of each layer is obtained by 2D convolutions encoding and decoding. The estimated depth map of the previous layer is upsampled as the reference depth map of the next layer to reduce the number of depth samples in the next layer. Each layer performs similar encoding and decoding operations, The final depth map are estimated after three layers to achieve the coarse-to-fine effect. Our MVS depth estimation network is trained on the ScanNet dataset [28], with a scale-invariant loss function to accommodate the relative poses of the sparse visual odometers:
@@
where $g^{s}_{i,j}={\uparrow_{gt}\log\hat{D}^{s}_{i,j}-\log{D}^{gt}_{i,j}}$. $D^{gt}_{i,j}$ denotes a ground truth depth map, which is aligned to the size of predicted depth $D^{gt}_{i,j}$ by an upsampling operation $\uparrow_{gt}$. $\lambda$ is a constant 0.85.
In addition, the multi-view loss and the normal loss are added to the loss function to maintain geometric consistency of depth estimation:
@@
Differentiable Gaussian map representation. We use 3D Gaussian Splatting as a dense representation of the scene. 3D Gaussian Splatting is an efficient 3D spatial representation that maps scenes in a way similar to point clouds, avoiding the need to sample in the blank areas of the space like NeRF and having the ability to render in real-time. The influence of a single 3D Gaussian $p_{i}\in\mathbb{R}^{3}$ in 3D scene is as follows:
@@
where $\pi(\cdot)$ denotes the projection of the 3D Gaussian center, $\textbf{T}_{CW}\in SE(3)$ is the the transformation matrix from world coordinate to camera coordinate in 3D space, $J$ is a linear approximation to the Jacobian matrix of the projective transformation, $W$ is the rotational component of $\textbf{T}_{CW}$. The Eq. (8), (9) is differentiable, which ensures that the Gaussian map can be used with first-order gradient descent to continuously optimize the geometric and photometric of the map, allowing the map to be rendered as photo-realistic images. A single pixel color $C_{p}$ is rendered from $N$ Gaussians by splatting and blending:
@@
Figure 4: Depth smooth regularization loss. Comparing the effect of having no depth smooth loss, there is better photometry and geometry in the left with depth smooth loss, and bad photometry and geometry in the right.
Mapping Optimization Losses. We changed the loss function of the original 3D Gaussian splatting and added more geometric constraints to make it more suitable for online mapping systems like SLAM. Specifically, our losses function consists of four components: photometric loss, depth geometric loss, depth smooth regularization loss and isotropic loss. In the photometric loss, the L1 loss is calculated between the rendered color image and the ground truth color image in the current camera pose $\textbf{T}_{CW}$:
$\mathcal{L}_{pho}=\|I(\mathcal{G},\textbf{T}_{CW})-I^{gt}\|_{1}$ (11)
where $I(\mathcal{G},\textbf{T}_{CW})$ is the rendered color image from Gaussians $\mathcal{G}$, and $I^{gt}$ is ground truth color image.
@@
The prior depth maps obtained from the MVS network may not be entirely accurate. As depicted in Fig. 4, direct utilization of this depth maps lead to erroneous guidance in the geometric reconstruction of the Gaussian map. Therefore, we introduce the depth smooth regularization loss:
@@
where $d_{i,j}$ denotes the depth value of the pixel coordinate at $(i,j)$ in the rendering depth map, it reduces the effect of errors in priori depth maps by regularizing neighboring pixels in rendered depth maps.
The original 3D Gaussian Splatting algorithm places no constraints on the Gaussians in the ray direction along the viewpoint. This has no effect on 3D reconstruction with fixed viewpoints. However, SLAM is an online mapping system, so this causes the Gaussians to elongate along the direction of the view ray, leading to the appearance of artifacts. To solve this problem, as well as [23], we also introduce isotropic loss:
@@
Camera poses optimization from the Gaussian map. We use the camera pose $\textbf{T}^{S}_{CW}$ obtained from sparse visual odometry tracking in the frontend as the initial pose for Gaussian mapping in the backend. As in Eq. (10) and Eq. (12), we render the color image and depth map from the Gaussian map at the viewpoint of the current initial pose and compute the losses of renderings and the ground truth. Since this process is differentiable, the loss gradient is propagated to both the Gaussian map and the initial pose during the gradient backward process. The equation of the initial pose optimization update is as follows:
$\underset{\textbf{T}^{S}_{CW},\mathcal{G}}{\mathrm{arg\,min}}\,\mathcal{L}_{mapping}(\mathcal{G},\textbf{T}^{S}_{CW},I^{gt},\bar{D}_{d})$ (17)
where $\mathcal{L}_{mapping}$ is the Eq. (16), $I^{gt}$ and $\bar{D}_{d}$ is ground truth color image and optimized priori depth map from the viewpoint of $\textbf{T}^{S}_{CW}$. Minimize the mapping loss to optimize both Gaussians $\mathcal{G}$ and initial pose $\textbf{T}^{S}_{CW}$ simultaneously.

### v2

@@ Abstract
This letter introduces a novel framework for dense Visual Simultaneous Localization and Mapping (VSLAM) based on Gaussian Splatting. Recently, SLAM based on Gaussian Splatting has shown promising results. However, in monocular scenarios, the Gaussian maps reconstructed lack geometric accuracy and exhibit weaker tracking capability. To address these limitations, we jointly optimize sparse visual odometry tracking and 3D Gaussian Splatting scene representation for the first time. We obtain depth maps on visual odometry keyframe windows using a fast Multi-View Stereo (MVS) network for the geometric supervision of Gaussian maps. Furthermore, we propose a depth smooth loss and Sparse-Dense Adjustment Ring (SDAR) to reduce the negative effect of estimated depth maps and preserve the consistency in scale between the visual odometry and Gaussian maps. We have evaluated our system across various synthetic and real-world datasets. The accuracy of our pose estimation surpasses existing methods and achieves state-of-the-art. Additionally, it outperforms previous monocular methods in terms of novel view synthesis and geometric reconstruction fidelities.
Index Terms:
SLAM; Mapping; 3D Gaussian Splatting
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a key technology in robotics and autonomous driving. It aims to solve the problem of how robots determine their location and reconstruct maps of the environment in unknown scenes. The development of SLAM technology has gone through multiple stages, starting with the initial filter-based method [1], advancing to graph optimization-based method [2], and more recently, integrating deep learning. This integration has significantly improved the accuracy and robustness of SLAM systems. With the rapid development of deep learning technology, a new approach to SLAM technology has emerged, utilizing differentiable rendering. The initial applications of differentiable rendering-based SLAM utilized Neural Radiance Fields (NeRF) as their foundational construction method. NeRF, as detailed in [3], employs neural networks to represent 3D scenes, enabling the synthesis of high-quality images and the recovery of dense geometric structures from multiple views. NeRF-based SLAM systems preserve detailed scene information during mapping, which enhances support for subsequent navigation and path planning. However, NeRF’s approach requires multiple forward predictions for each pixel during image rendering, leading to significant computational redundancy. Consequently, this inefficiency prevents NeRF-based SLAM from operating in real-time, thus limiting its practicality for immediate downstream tasks.
@@
Recently, a novel scene representation framework called 3D Gaussian Splatting [4] has demonstrated superior performance compared to NeRF. It features a more concise scene representation method and real-time rendering capability. This method not only delivers an accurate description of the scene but also offers a differentiable approach for optimizing the scene and camera poses. This opens up a new research direction for differentiable rendering-based SLAM. However, current Gaussian Splatting-based SLAM systems rely on the depth maps input to achieve precise geometric reconstruction, which constrains the scope of their application.
Figure 2: System pipeline. The system inputs an RGB stream and operates frontend and backend processes in parallel. In the frontend, sparse visual odometry extracts patch features from images to estimate poses. These estimated poses and images are inputs to a pre-trained Multi-View Stereo (MVS) network, which estimates priori depth maps. In the backend, the estimated priori depth maps and images, coupled with poses from the frontend, are utilized as supervisory information to construct a Gaussian map. The frontend and backend maintain scale consistency through the SDAR strategy.
This letter presents MGS-SLAM, a novel monocular Gaussian Splatting-based SLAM system. This work introduces several groundbreaking advancements in the field of SLAM, which include integrating Gaussian Splatting techniques with sparse visual odometry, employing a pre-trained Multi-View Stereo (MVS) depth estimation network, pioneering a geometric smooth depth loss, and developing the SDAR strategy to ensure scale consistency. Together, these innovations significantly improve the accuracy and functionality of SLAM systems that rely solely on RGB image input. Fig. 1 illustrates the map construction process: initially, sparse visual odometry constructs the sparse maps; subsequently, the MVS depth estimation network generates priori depth maps; these depth maps, along with the sparse point maps, are then refined through depth optimization in the SDAR; and finally, the Gaussian map is constructed using the optimized depth maps and depth smooth regularization loss.
@@
  • Introducing the first SLAM system that jointly optimizes sparse visual odometry poses and 3D Gaussian Splatting to achieve the accurate geometric reconstruction of Gaussian maps and pose tracking.
  • Developing a pre-trained Multi-View Stereo (MVS) depth estimation network that utilizes sparse odometry keyframes and their poses to estimate prior depth maps, thus providing crucial geometric constraints for Gaussian map reconstruction with only RGB image input.
@@ II Related Works
Monocular Dense SLAM. Over the past few decades, monocular dense SLAM technology has seen significant advancements. DTAM [5] pioneered one of the earliest real-time dense SLAM systems by performing parallel depth computations on GPU. To balance computational costs and accuracy, there are also semi-dense methods such as [6], but these methods struggle to capture areas with poor texture. In the era of deep learning, DROID-SLAM [7] utilizes optical flow networks to establish dense pixel correspondences and achieve precise pose estimation. Another study [8], combines a real-time VO system with a Multi-View Stereo (MVS) network for parallel tracking and dense depth estimation, and then the Truncated Signed Distance Function (TSDF) is used to fuse depth maps and extract mesh. Codemapping [9] and Rosinol et al. [10] incorporate sparse point cloud correction and volumetric fusion strategy on the estimated depth map to mitigate the impact of errors in the estimated depth map. We have also adopted a strategy for correcting the estimated depth map, but the difference is that we use a linear variance correction as depth optimization, which is less computations.
@@
Differentiable Rendering SLAM. With the emergence of Neural Radiance Fields (NeRF) in 2020, numerous NeRF-based SLAM works have been proposed. iMAP [11] represented the pioneering work in NeRF-based SLAM, utilizing a dual-threading mode to track camera poses and execute mapping simultaneously. NICE-SLAM [12] introduced feature grids based on iMAP, enabling NeRF-based SLAM to represent larger scenes. Subsequent works such as GO-SLAM [13] and Loopy-SLAM [14] incorporated global bundle adjustment (BA) and loop closure correction, further enhancing pose estimation accuracy and mapping performance. PLGSLAM [15] proposes a progressive scene representation method to improve reconstruction and localization accuracy in large scenarios. Recently, 3D Gaussian Splatting has shown superior performance in 3D scene representation. It has fast rendering capability and is more suitable for online systems like SLAM. SplaTAM [16] and GS-SLAM [17] combine 3D Gaussian Splatting with SLAM, leveraging the realistic scene reconstruction ability of 3D Gaussian Splatting to surpass NeRF-based SLAM methods in rendering quality. Compact-SLAM [18] proposes a compact 3D Gaussian Splatting SLAM system that reduces the number and the parameter size of Gaussian ellipsoids. NGM-SLAM [19] utilizes neural radiance field submaps for progressive scene expression, achieving effective loop closure detection. MonoGS [20] and Photo-SLAM [21] achieve monocular map reconstruction of Gaussian Splatting-based SLAM. However, existing Gaussian Splatting-based SLAM implementations typically require depth map input from RGB-D sensors to obtain accurate geometry reconstruction.
@@
Figure 3: The fast Multi-View Stereo network. The inputs of the network are images with poses from sparse visual odometry, image features are extracted by Feature Pyramid Network (FPN) and warped to the 2D cost volume. Finally, encoded and decoded to depth maps using coarse-to-fine strategy.
@@ III Methods
Our approach utilizes RGB image as input, parallelly performing camera pose estimation and photorealistic dense mapping. As depicted in Fig. 2, the core idea of the approach is to use a pre-trained Multi-View Stereo (MVS) network to couple sparse VO and dense Gaussian Splatting mapping. Specifically, in the frontend part, tracking RGB image provides the backend with coarse camera poses and priori depth maps (Sec. III-A). In the backend part, we represent the dense map using 3D Gaussian Splatting, and jointly optimize the dense map and the coarse poses from the frontend (Sec. III-B). In the system components part, system initialization, selecting the keyframes for the system and correcting the scale between sparse point cloud map and dense Gaussian map by SDAR strategy are reported (Sec. III-C).
@@ III-A Sparse Visual Odometry Frontend
To achieve more accurate camera pose tracking and provide dense depth geometry before backend mapping, the frontend of our framework is built on the Deep Patch Visual Odometry (DPVO) [22] algorithm. DPVO is a learning-based sparse monocular visual odometry method. Given an input RGB stream, the scene is represented as a collection of camera poses $\textbf{T}\in SE(3)^{N}$ and a series of square image patches P extracted from the images. The reprojection of a square patch $k$ taken from frame $i$ in frame $j$ can be formulated as:
@@
The core of DPVO is an update operator that computes the hidden state for each edge $(k,i,j)\in\varepsilon$. It optimizes the reprojection errors on the patch graph to predict a 2D correction vector $\delta^{ij}_{k}\in\mathbb{R}^{2}$ and confidence weight $\psi^{ij}_{k}\in\mathbb{R}^{2}$. Bundle Adjustment (BA) is performed using optical flow correction as a constraint, with iterative updates and refinement of camera poses and patch depths achieved through the non-linear least squares method. The cost function for bundle adjustment is as follows:
@@
Multi-view priori depth estimation. The backend dense Gaussian mapping requires the geometric supervision of depth maps to obtain the accurate geometric positions of Gaussians. In order to make monocular SLAM have the ability of geometric supervision, unlike the previous method [23], we use a pre-trained Multi-View Stereo (MVS) network to estimate priori depth maps on the keyframes window of DPVO, the network is shown in Fig. 3. This method utilizes the geometric consistency of the MVS network to achieve the supervision of the geometric positions of Gaussians through only monocular RGB image input. Furthermore, our MVS network consists entirely of 2D convolutions with a coarse-to-fine structure that progressively refines the estimated priori depth map to reduce the runtime of the MVS network. Tab. IV-B and Tab. V show that this method achieves better rendering and reconstruction performance.
To be more specific, the frame currently tracked by the sparse visual odometry is used as the reference image $\textbf{I}^{0}$. Additionally, we employ the previous N keyframes as a series of original images $\textbf{I}^{n\in{1,...,N}}$. These images and their corresponding camera poses, serve as inputs to the MVS network. Utilizing the Feature Pyramid Network (FPN) module, we extract three layers of image features $\textbf{F}^{s}_{i}$ for each image, with $s$ denoting the layer index and $i$ representing the image index. In each layer, the original image features dot the reference image features by a differentiable warping operation to obtain a cost volume with dimensions $\textbf{D}\times\textbf{H}^{s}\times\textbf{W}^{s}$, and the priori depth map of each layer is obtained by 2D convolutions encoding and decoding. The estimated depth map of the previous layer is upsampled as the reference depth map of the next layer. The final depth map is estimated after three layers to achieve the coarse-to-fine effect.
Our MVS depth estimation network is trained on the ScanNet dataset [24]. We train with the AdamW optimizer for 100k steps with a weight decay of $10^{-4}$, and a learning rate of $10^{-4}$ for 70k steps, $10^{-5}$ until 80k, then dropped to $10^{-6}$ for remainder, which takes approximately 84 hours on two 24GB RTX3090 GPUs. We use a scale-invariant loss function to accommodate the relative poses of the sparse visual odometers:
@@
where $g^{s}_{i,j}={\uparrow_{gt}\log\hat{D}^{s}_{i,j}-\log{D}^{gt}_{i,j}}$. $D^{gt}_{i,j}$ denotes a ground truth depth map, which is aligned to the size of predicted depth $D^{s}_{i,j}$ by an upsampling operation $\uparrow_{gt}$. $\lambda$ is a constant 0.85.
In addition, the multi-view loss and the normal loss are added to the loss function to maintain the geometric consistency of depth estimation. The multi-view loss average absolute error on log depth over all valid points:
@@
Differentiable Gaussian map representation. We use 3D Gaussian Splatting as a dense representation of the scene. The influence of a single 3D Gaussian $p_{i}\in\mathbb{R}^{3}$ in 3D scene is as follows:
@@
where $\pi(\cdot)$ denotes the projection of the 3D Gaussian center, $\textbf{T}_{CW}\in SE(3)$ is the the transformation matrix from world coordinate to camera coordinate in 3D space, $J$ is a linear approximation to the Jacobian matrix of the projective transformation, $W$ is the rotational component of $\textbf{T}_{CW}$. The Eq. (8) and Eq. (9) are differentiable, which ensures that the Gaussian map can be used with first-order gradient descent to continuously optimize the geometric and photometric of the map, allowing the map to be rendered as photo-realistic images. A single pixel color $C_{p}$ is rendered from $N$ Gaussians by splatting and blending:
@@
Figure 4: Depth smooth regularization loss. Comparing the effect of having no depth smooth loss, there is better photometry and geometry with depth smooth loss, and bad photometry and geometry without depth smooth loss.
Mapping Optimization Losses. We changed the loss function of the vanilla 3D Gaussian splatting and added more geometric constraints to make it more suitable for online mapping systems like SLAM. Specifically, our loss function consists of four components: photometric loss, depth geometric loss, depth smooth regularization loss and isotropic loss. In the photometric loss, the L1 loss is calculated between the rendered color image and the ground truth color image in the current camera pose $\textbf{T}_{CW}$:
$\mathcal{L}_{pho}=\|I(\mathcal{G},\textbf{T}_{CW})-I_{gt}\|_{1}$ (11)
where $I(\mathcal{G},\textbf{T}_{CW})$ is the rendered color image from Gaussians $\mathcal{G}$, and $I_{gt}$ is ground truth color image.
@@
The prior depth maps obtained from the MVS network may not be entirely accurate. As depicted in Fig. 4, direct utilization of these depth maps leads to erroneous guidance in the geometric reconstruction of the Gaussian map. Similar to NeSLAM [25], we introduce the depth smooth regularization loss to reduce this erroneous guidance:
@@
where $d_{i,j}$ denotes the depth value of the pixel coordinate at $(i,j)$ in the rendering depth map. However, NeSLAM is an RGB-D SLAM system, which optimizes the noisy depth from RGB-D sensors by the denoising network and constraining the standard variance of depth to obtain better depth input. In contrast, we regularize the adjacent pixels between depth maps rendered from the Gaussian map, enabling the Gaussians to have better geometric positions.
The vanilla 3D Gaussian Splatting algorithm places no constraints on the Gaussians in the ray direction along the viewpoint. This has no effect on 3D reconstruction with fixed viewpoints. However, SLAM is an online mapping system, so this causes the Gaussians to elongate along the direction of the view ray, leading to the appearance of artifacts. To solve this problem, as well as [20], we also introduce isotropic loss:
@@
Camera poses optimization from the Gaussian map. We use the camera poses $\textbf{T}^{i}_{CW}$ obtained from sparse visual odometry tracking in the frontend as the initial poses for Gaussian mapping in the backend. As in Eq. (10) and Eq. (12), we render the color image and depth map from the Gaussian map at the viewpoint of the current initial poses and compute the loss of renderings and the ground truth. Since this process is differentiable, the loss gradient is propagated to both the Gaussian map and the initial poses during the gradient backward process. The equation of the initial poses optimization update is as follows:
$\underset{\textbf{T}^{i}_{CW},\mathcal{G}}{\mathrm{arg\,min}}\,\sum_{i=1}^{n}\mathcal{L}_{mapping}(\mathcal{G},\textbf{T}^{i}_{CW},I^{i}_{gt},\bar{D}^{i}_{d})$ (17)
where $\mathcal{L}_{mapping}$ is the Eq. (16), $I^{i}_{gt}$ and $\bar{D}^{i}_{d}$ are $i$th ground truth color image and optimized priori depth map from the viewpoint of $\textbf{T}^{i}_{CW}$ in mapping window. $n$ is mapping window size. Minimize the mapping loss to optimize both Gaussians $\mathcal{G}$ and initial poses $\textbf{T}^{i}_{CW}$ simultaneously.